Metallica Bearings, Inc., is a young start-up company. No dividends will be paid on the stock over the next nine years because the firm needs to plow back its earnings to fuel growth. The company will pay a dividend of $17 per share 10 years from today and will increase the dividend by 3.9 percent per year thereafter. If the required return on this stock is 12.5 percent, what is the current share price? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

g= 3.9% 17. EX.039 + + Jo 1 Dio=17 All = 17+ (1+g) = 17*(1.039) 17.66 Terminal value (Pio) = Re-g 17X 1.039 - Re= 12.5% 3 •12